Using the sample ina family of four children, construct a probability distribution for the variable Z representing the number of boys. Draw the histogram of the probability distribution.

1
Expert's answer
2022-02-13T17:46:39-0500

There are "2^4=16" possible outcomes.


"P(Z=0)=P(GGGG)=(\dfrac{1}{2})^4=\dfrac{1}{16}"

"P(Z=1)=P(GGGB)+P(GGBG)"

"+P(GBGG)+P(BGGG)=4(\dfrac{1}{2})^4=\dfrac{1}{4}"

"P(Z=2)=P(GGBB)+P(GBBG)"

"+P(GBGB)+P(BGGB)+P(BGBG)"

"+P(BBGG)=6(\dfrac{1}{2})^4=\dfrac{3}{8}"

"P(Z=3)=P(GBBB)+P(BGBB)"

"+P(BBGB)+P(BBBG)=4(\dfrac{1}{2})^4=\dfrac{1}{4}"

"P(Z=4)=P(BBBB)=(\dfrac{1}{2})^4=\dfrac{1}{16}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 z & 0 & 1 & 2 & 3 & 4 \\ \hline
 p(z) & 1/16 & 1/4 & 3/8 & 1/4 & 1/16 \\
 
\end{array}"